Multiply mixed numbers 1 2/3 × 2 5/6
1
Expert's answer
2019-11-19T10:15:12-0500

Solution. Convert fractions and multiply


"1\frac {2}{3} \times 2 \frac {5} {6}=\frac {5}{3}\times\frac {17}{6}=\frac {85}{18}=4 \frac {13}{18}"

Answer.


"\frac {85}{18}=4 \frac {13}{18}"